When 2.35g of Mg(OH)2 is added to 250.0 mL of water, the temperature of the water raises from 20.5°C to 36.0°C. Calculate the molar enthalpy of solution.
1
Expert's answer
2020-09-17T07:18:42-0400

M(Mg(OH)2) = 24.3 +16*2 + 2 = 58.3

n(Mg(OH)2) = 2.35 / 58.3 = 0.04 mol

Q = Cp(H2O) * (T2 - T1)*m(H2O) =  4179.6 * (36 - 20.5) * 0.25 = 16195.95 J

Hsolution= 16195.95 / 0.04 = 404898.75 J/mol = 404.9 kJ/mol